[Hallc_running] [New Logentry] RC Report - 8/25/22

arunts at jlab.org arunts at jlab.org
Thu Aug 25 10:20:02 EDT 2022

Logentry Text:
We finished taking LH2, 3/4 and dummy runs (page 33 of the run plan) during swing and early owl shifts. We then moved on to Heep #4 setting (page 16), took data and then on to Heep #5 setting (page 18). We were briefly limited to 75 uA due to BCM temperature regulation issues (https://logbooks.jlab.org/entry/4033600) and were taken to controlled access mode to debug this issue. We are now back in business at 80 uA. We started to get quite a few alarms from SHMS preshower13 even after a quick power cycle of this channel. So Dave Gaskell moved this channel from crate 11, slot, channel 12 to crate 4, slot 3, channel 14, updated the HV.hvc file and recompiled the gui. The alarms have gone for now but we will continue to monitor this channel. The plan is to move on to negative polarity runs (page 40 of the run plan) where we will stay for ~40 hours. 

Let's have a brief RC meeting today at 11:00 AM EST.

Meeting ID: 160 710 0885 - Passcode: 595071


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4033679
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220825/202e11f4/attachment.html>

More information about the Hallc_running mailing list